Sage Corporation had a net income for the current year ending December 31, 2021 of $1,248,000. Throughout 2021 the following items were outstanding:

389,500 common shares
18,000 Class A $3 cumulative preferred shares that were convertible to common shares at a rate of 1:1
49,500 Class B $4 non-cumulative preferred shares that were convertible at a rate of one common share for every two preferred shares.
$496,000, 8% bonds that were convertible to 16,000 common shares
$295,000, 10% bonds convertible to 10,000 common shares

No dividends were declared or paid in 2021. Sage’s tax rate is 25%.

1. Calculate the income effect of the dividends on Class A preferred shares.

2. Calculate the income effect of the dividends on Class B preferred shares.

3. Calculate basic earnings per share. (For simplicity, ignore the requirement to record the debt and equity portions of the convertible bond separately).

4. Calculate the after-tax interest paid on the 8% bonds.

After-tax interest on 8% bonds converted: $ _________

5. Determine an incremental per share effect for 8% bonds.

Potentially dilutive security
Incremental
Numerator Effect
Incremental
Denominator Effect
EPS
8% Bonds $ $ $

6. Calculate the after-tax interest paid on the 10% bonds.

After-tax interest on 10% bonds converted: $____________

7. Determine an incremental per share effect for 10% bonds.

Potentially dilutive security
Incremental
Numerator Effect
Incremental
Denominator Effect
EPS
10% Bonds $ $ $

8. Determine an incremental per share effect for Class A preferred shares.

Potentially dilutive security
Incremental
Numerator Effect
Incremental
Denominator Effect
EPS
Class A preferred shares $ $ $

9. Determine an incremental per share effect for Class B preferred shares.

Potentially dilutive security
Incremental
Numerator Effect
Incremental
Denominator Effect
EPS
Class B Preferred shares $ $ $

9. Rank the potentially dilutive securities from most dilutive to least dilutive. (Options: Rank 1, Rank 2, Rank 3, Rank 4, Anti-dilutive)

8% bonds
10% bonds
Class A preferred shares
Class B preferred shares

10. Calculate diluted earnings per share.

Numerator Denominator EPS
Basic $ $ $
$ $
Sub Total $ $ $
$ $
Sub Total $ $ $
$ $
$ $ $

Diluted EPS: __________
Answer #1

Total No of Common shares:

1 Common shares (given) 389500
2 Converted Class A @1:1 18000
3 Converted Class B @1:2 24750
4 8% bonds 16000
5 10% Bonds 10000
Total Common Shares 458250

Total Earnings : $12,48,000

Basic Earnings Per Share: $1248000/458250 = $2.72 (Answer for Qn 3)

Net Income $12,48,000
Tax on Earnings :$1248000*25%= $3,12,000
After-tax available earnings for distribution of Dividends = $9,36,000

Answer for Qn4 :after-tax interest paid on the 8% bonds.

After-tax interest on 8% bonds converted: $ 496000*8%=$39,680

Answer for Qn6: after-tax interest paid on the 10% bonds.

After-tax interest on 10% bonds converted: $ 295000*10%=$29,500